# TokMem: Tokenized Procedural Memory for Large Language Models

### v1

@@ Abstract
Large language models rely heavily on prompts to specify tasks, recall knowledge and guide reasoning. However, this reliance is inefficient as prompts must be re-read at each step, scale poorly across tasks, and lack mechanisms for modular reuse. We introduce TokMem, a tokenized procedural memory that stores recurring procedures as compact, trainable embeddings. Each memory token encodes both an address to a procedure and a control signal that steers generation, enabling targeted behavior with constant-size overhead. To support continual adaptation, TokMem keeps the backbone model frozen, allowing new procedures to be added without interfering with existing ones. We evaluate TokMem on 1,000 tasks for atomic recall, and on function-calling tasks for compositional recall, where it consistently outperforms retrieval-augmented generation while avoiding repeated context overhead, and fine-tuning with far fewer parameters. These results establish TokMem as a scalable and modular alternative to prompt engineering and fine-tuning, offering an explicit procedural memory for LLMs.
@@ 1 Introduction
Large language models (LLMs) have become the foundation of modern natural language processing, powering a wide range of applications in text understanding, generation, and coding (Brown et al., 2020; Grattafiori et al., 2024). Prompting is a widely adopted way to steer LLM behavior, where in-context learning enables adaptation to new tasks without parameter updates (Brown et al., 2020). Consequently, prompt and context engineering has emerged as a dominant mechanism for specifying tasks, obtaining relevant information, and guiding multi-step reasoning or tool invocation (Wei et al., 2022b; Yao et al., 2023; Sahoo et al., 2025).
Despite its success, this reliance on long prompts is inherently inefficient. Constructing and maintaining prompts are labor-intensive and difficult to scale across many tasks (Liu et al., 2023). At inference, long prompts increase computational cost because the attention mechanism scales quadratically with sequence length (Vaswani et al., 2017), and they reduce the effective context window available for inputs and outputs, often leading to truncation and loss of details (Liu et al., 2024a). These limitations make it difficult to manage expanding tasks and to execute procedures efficiently.
@@
To address these issues, recent approaches offload prompts into retrieval-based memory. Retrieval-augmented generation (RAG) (Lewis et al., 2020) and memory systems such as MemGPT(Packer et al., 2023) fetch and reinsert documents or conversational state at inference time. While retrieving in-context learning demonstrations (Wei et al., 2022b) can provide procedural cues that guide the model’s behavior, the mechanism still largely aligns with declarative memory in cognitive science: knowledge remains as explicit text that must be repeatedly interpreted. This creates two challenges: (1) retrieved content still occupies the context window, reintroducing quadratic compute and truncation pressure, and (2) frequently used procedures are repeatedly re-read as text rather than compiled into compact, reusable procedures, missing the compression opportunity suggested by minimum description length principles (Grünwald, 2007).
@@
We propose Tokenized Memory (TokMem), an explicit form of procedural memory that encodes recurring procedures as compact, trainable tokens while keeping the backbone frozen. Each memory token serves both as an address to a procedure and as a control signal that steer generation, enabling targeted behavior with constant-size overhead. Specifically, rather than front-loading procedures as long prompts, TokMem integrates memory tokens directly into the generation process. As shown in Figure 1b, memory tokens can be invoked and chained across stages: after producing one response segment, the model retrieves the next relevant token, which conditions the next stage, enabling the composition of multi-step behaviors such as parsing, searching, and formatting.
@@
A key advantage of TokMem is that its memory tokens are parameter-isolated from the backbone. This design ensures that the learned procedural knowledge is fully stored in dedicated tokens, allowing new procedures to be added without interfering with existing ones. TokMem thus naturally supports continual learning, where the model can accumulate procedural skills over time while preserving stability. This capability mirrors human procedural memory, where skills are gradually acquired through practice and later invoked by contextual cues (Anderson & Lebiere, 1998). In this way, TokMem enables both efficient learning and continual expansion of procedural knowledge.
@@
We evaluate TokMem in two complementary settings. In the atomic memory recall setting, each task from Super-Natural Instructions (Wang et al., 2022a) is treated as a distinct procedure, TokMem stores and retrieves 1,000 such procedures efficiently, without catastrophic forgetting. In the compositional memory recall setting based on function-calling tasks (Liu et al., 2024b), each tool invocation is modeled as an atomic procedure, and solving a query requires chaining multiple procedures together. TokMem supports this process by composing memory tokens, enabling the model to assemble procedures into coherent multi-step behaviors. Across both settings and multiple LLM backbones, TokMem consistently outperforms retrieval-based baselines and surpasses parametric fine-tuning while using far fewer trainable parameters.
@@
Figure 1: Overview of TokMem. (a) New memory (colored) tokens are interleaved with text sequences, learning with next-token-prediction while the LLM backbone remains frozen. (b) An example of inference, a query recalls and chains memory tokens (parse, search, format), enabling multi-step procedural behavior without long prompts.
@@ 2 Method
We begin by reviewing how Transformer-based LLMs process input sequences and then describe how TokMem departs from existing approaches to enable procedural memory.
@@ 2.1 Textualized Context Engineering
A Transformer (Vaswani et al., 2017) processes a sequence of tokens $(a_{1},\dots,a_{n})\in{\mathbb{N}}^{n}$, where each $a_{i}$ is an integer representing the index of a token (usually sub-words). The model retrieves the corresponding embedding vector from the embedding layer and produces an input sequence $({\bm{x}}_{1},\dots,{\bm{x}}_{n})\in{\mathbb{R}}^{n\times d}$, which is then consumed to predict the next token in sequence.
Recent advances in prompting can be viewed as textualized context engineering, where the goal is to carefully choose input tokens that steer the model toward improved behavior. For example, chain-of-thought prompting (Wei et al., 2022a) augments input with intermediate reasoning steps to strengthen logical inference. Retrieval-based methods such as RAG (Lewis et al., 2020) and memory-augmented approaches like MemGPT (Packer et al., 2023) provide relevant information in text form by retrieving external memory.
Concretely, to evoke a procedural response $(r_{1},\ldots,r_{n})$ for a query $(q_{1},\ldots,q_{k})$, these methods may prepend textual context $(c_{1},\ldots,c_{m})$ that describes the procedure:
$(c_{1},\ldots,c_{m},\;\;q_{1},\ldots,q_{k})\;\xrightarrow{\text{LLM}}\;(r_{1},\ldots,r_{n}),\qquad m\gg k.$ (1)
While effective, these approaches are costly: the same procedural context must be re-read at every invocation, which can exhaust the limited context window (Agarwal et al., 2024).
@@ 2.2 TokMem: procedural memory as a token
Our key idea is that frequently reused procedures can be effectively “compressed” and stored by encoding them into an internalized memory token, bypassing repeated textual specification. Specifically, we create a memory bank of $l$ special embeddings:
@@
Each ${\bm{m}}_{i}$ is a trainable vector with no direct textual translation and represents a unique procedure. For simplicity, we label each memory token ${\bm{m}}_{i}$ to have a special index $a_{m_{i}}\in{\mathbb{N}}$. In contrast to textualized context engineering, TokMem evokes a procedural response without any explicit context:
$(q_{1},\ldots,q_{k})\;\xrightarrow{\text{LLM + }M}\;(r_{1},\ldots,r_{n}),$ (3)
where the model recalls the appropriate memory token $a_{m_{i}}$ internally based on the query. This eliminates redundant context engineering and achieves $O(1)$ procedural invocation.
To connect these tokens with training, we define a procedure–response pair as the concatenation of a memory token $a_{m_{i}}$ and its corresponding response tokens $(r_{i1},\dots,r_{in})$. Importantly, each training instance may contain multiple such pairs to model multi-stage reasoning or composition (Khot et al., 2023; Zhou et al., 2023). Formally, the sequentialized training sequence has the layout
${\bm{a}}=\big(q_{1},\ldots,q_{k},\;\;\underbrace{a_{m_{i}},a_{r_{i1}},a_{r_{i2}},\ldots}_{\text{procedure-response pair}},\;\;\underbrace{a_{m_{j}},a_{r_{j1}},a_{r_{j2}},\ldots}_{\text{procedure-response pair}},\;\;\ldots\big).$ (4)
@@
$\mathcal{L}({\bm{a}};M)\;=\;-\sum_{i>k}\log\Pr\big(a_{i}\;\big|\;{\bm{a}}_{<i};M\big).$ (5)
During optimization, only the memory embeddings $({\bm{m}}_{1},\dots,{\bm{m}}_{l})$ are updated, while the Transformer backbone remains frozen. We visualize our training process in Figure 1a.
@@ 2.3 Stabilizing New Memories
TokMem allows new procedures to be added incrementally to the memory bank, mimicking how humans continually form new procedural memories without disrupting existing skills (Anderson & Lebiere, 1998; Squire, 2009). This design enables practical deployment scenarios, where an LLM can steadily accumulate routines across domains and tasks rather than retraining from scratch.
However, adding new tokens poses stability challenges. If all procedural memories are introduced at once, the model risks overfitting to spurious patterns. Conversely, when new embeddings are added gradually, they often develop inflated norms that dominate routing logits and suppress older memories. To address this, we introduce renormalization, which is a lightweight post-update calibration to the memory bank $M\in{\mathbb{R}}^{l\times d}$.
Let $A$ and $I$ denote the indices of the active (new) and inactive (existing) procedural memories, respectively. We estimate the prevailing scale from the inactive set:
$\bar{n}_{I}\;=\;\frac{1}{|I|}\sum_{j\in I}\|{\bm{m}}_{j}\|_{2},$ (6)
and rescale each active embedding as
${\bm{m}}_{i}\;\leftarrow\;{\bm{m}}_{i}\cdot\frac{\bar{n}_{I}}{\|{\bm{m}}_{i}\|_{2}+\varepsilon},\qquad\forall\,i\in A.$ (7)
This operation preserves the directions of newly added embeddings while aligning their magnitudes to the established scale of the memory bank, ensuring smooth integration without overwhelming the routing dynamics. The computational overhead is negligible, scaling as $O(|A|d)$.
@@ 3 Experiments
We evaluate TokMem in two complementary scenarios that test scalability and compositionality:
  • Atomic Memory Recall: Each task from the Super-Natural Instructions dataset (Wang et al., 2022a) is framed as a standalone procedure, where a query directly maps to the desired response. This tests TokMem’s ability to store many independent procedures without catastrophic forgetting.
  • Compositional Memory Recall: Evaluated on the function-calling dataset (Liu et al., 2024b), where invoking a tool is treated as a procedure and solving a query requires composing several function calls. This tests whether learned memory tokens can be flexibly chained.
Experiments are conducted on Qwen (Qwen & et al., 2025) and Llama (Grattafiori et al., 2024) model families, ranging from the 0.5B-parameter Qwen to the 8B-parameter Llama model.
@@ 3.1 Experimental Setup
Across both settings, we compare TokMem with textualized context engineering, retrieval-augmented memory, and parameter-efficient fine-tuning.
  • Base: In the atomic setting, the model answers queries without demonstrations, providing a non-parametric lower bound highlighting the need to recall task knowledge.
  • ICL: In the compositional setting, we augment input with all tool descriptions and prepend two compositional procedure–response demonstrations, representing a context engineering baseline.
  • RAG: We retrieve relevant demonstrations or tool usages with Sentence-BERT (Reimers & Gurevych, 2019) and prepend them to the query, following memory-augmented generation (Packer et al., 2023; Chhikara et al., 2025; Xu et al., 2025).
  • Fine-tuning: We use low-rank adapters (Hu et al., 2022) inserted into the query and key projections of the transformer, updating millions of parameters depending on model sizes. This serves as a parametric form of procedural memory, but is prone to forgetting as new tasks are introduced.
  • Replay Memory: To mitigate catastrophic forgetting in fine-tuning, we follow the idea of experience replay (Mnih et al., 2015) by maintaining a buffer of previously seen precedures and mixing them with the current training data.
@@
All methods are implemented in HuggingFace Transformers and trained on a single NVIDIA A6000 GPU with 48GB memory using mixed-precision (bfloat16) training. The backbone models remain frozen; for fine-tuning, only the adapter weights (rank $r=8$) are updated, while for TokMem, only the embeddings of the newly added procedure IDs are trainable. Tokenizer vocabulary is expanded with these procedure IDs, and their embeddings are initialized by averaging the pretrained embeddings (Hewitt, 2021). For Replay Memory, we mix $20\%$ of replayed samples with the current batch, using a buffer of $500$ examples refreshed every $10$ tasks in the atomic setting and 1,0001{,}000 examples updated each round in the compositional setting.
We optimize with AdamW using a learning rate $5\times 10^{-5}$ for fine-tuning and $5\times 10^{-3}$ for TokMem; weight decay is $10^{-2}$ for fine-tuning and zero for TokMem. Training runs for one epoch with batch size $4$ and maximum sequence length $1024$, using teacher forcing and applying the loss only to memory-token and response positions.
3.2 Atomic Memory Recall
Dataset Details.
We evaluate on the Super-Natural Instructions (SNI) dataset (Wang et al., 2022a), which provides diverse QA-style natural language tasks. Here, each task is treated as an individual procedure: a query directly invokes the learned procedure to produce the desired response. We sample 1,0001{,}000 English tasks, each task contains $500$ training and $50$ test examples. To reflect how memories are typically acquired over time, we introduce tasks sequentially during training rather than all at once. We scale the number of tasks from $10$ up to 1,0001{,}000, and record checkpoints after training on $\{10,50,200,500,1{,}000\}$ tasks. This resembles incremental domain adaptation (Asghar et al., 2020), where at each checkpoint, the performance is evaluated across all previously seen tasks. Additional tasks details are provided in Appendix D.1.
We follow Wang et al. (2022a) and use Rouge-L (Lin, 2004) to evaluate generation quality. For the methods with explicit memory routing (RAG and TokMem), we additionally report accuracy, reflecting whether the correct procedure was selected and applied.
Decoupled Memory Embeddings.
We also include an ablation variant where memory tokens are decoupled to an address token and a steering token. This decoupling separates the roles of a memory token and also increase the capacity of TokMem. We refer to this variant as TokMem with decoupled embeddings (TokMem+DC), with further details provided in Appendix A.
Table 1: Atomic recall performance on SNI, reported with Rouge-L. TokMem consistently outperforms fine-tuning and RAG across models and scales, maintaining strong results even at 1,000 tasks.
@@ 3.2 Atomic Memory Recall
Table 2: Task routing accuracy. TokMem achieves near-perfect routing accuracy at 1,000 tasks, far exceeding RAG retriever whose accuracy falls below 80%.

### v2

# TokMem: One-Token Procedural Memory for Large Language Models

@@ Abstract
Large language models are typically controlled via prompts, which must be repeatedly re-processed for every new query and are difficult to reuse modularly. We introduce TokMem, a procedural memory framework that compiles each reusable task procedure into a single trainable memory token. Each token serves as both a procedure index and a generation control signal that steers generation, enabling targeted behaviors with constant-size overhead. TokMem keeps the backbone LLM frozen and stores procedural knowledge entirely in these dedicated units, so new procedures can be added continually without interfering with existing ones. We evaluate TokMem on two settings: atomic recall over 1,000 Super-Natural Instructions tasks and compositional recall on multi-step function-calling. Our results show that TokMem consistently outperforms retrieval-augmented prompting while avoiding repeated context overhead. Moreover, it matches or exceeds parameter-efficient fine-tuning with substantially fewer trainable parameters. [^1] Our code is publicly available at https://github.com/MANGA-UOFA/TokMem
@@ 1 Introduction
Large language models (LLMs) have become the foundation of modern natural language processing, powering a wide range of applications in text understanding, generation, and coding (Brown et al., 2020; Llama Team, 2024; Chen et al., 2021). Prompting is a widely adopted way to steer LLM behavior, where in-context learning enables adaptation to new tasks without parameter updates (Brown et al., 2020). Consequently, prompt and context engineering has emerged as a dominant interface for specifying tasks, obtaining relevant information, and guiding multi-step reasoning or tool invocation (Wei et al., 2022b; Yao et al., 2023; Sahoo et al., 2025).
Despite their effectiveness, long prompts are inherently inefficient. They are labor-intensive to construct and maintain, and they do not scale well when conditioning diverse queries across different tasks (Liu et al., 2023). At inference time, longer contexts increase compute because self-attention scales quadratically (Vaswani et al., 2017). They also shrink the usable context window for inputs and outputs, often causing truncation and information loss (Liu et al., 2024a). Moreover, as tasks accumulate, these costs make it increasingly difficult to execute task-specific behaviors efficiently.
@@
To mitigate these issues, recent work externalizes prompts into retrieval-based memory. Retrieval-augmented generation (RAG; Lewis et al., 2020) is used to retrieves and reinserts documents or conversational state at inference time, a strategy exemplified by MemGPT (Packer et al., 2024). Although retrieving in-context demonstrations (Wei et al., 2022b) can supply procedural cues, the retrieved knowledge remains explicit text that must be repeatedly interpreted, resembling declarative memory in cognitive science. This leads to two drawbacks: (1) retrieved content still occupies the context window, reintroducing quadratic compute and truncation pressure, and (2) frequently used procedures are repeatedly re-read as text rather than distilled into compact, reusable representations.
We propose One-Token Procedural Memory (TokMem), a modular framework that encodes task procedures into compact, trainable tokens while maintaining a frozen LLM backbone. Here, a procedure refers to a reusable context–response mapping that captures a specific task behavior; the terminology is inspired by physiological research (Anderson and Lebiere, 1998), which posits procedural knowledge as a compiled form of skill that is more efficient than the slow interpretation of facts. Unlike factual knowledge that can be expressed in plain text, procedural knowledge (e.g., riding a bicycle) is often more nuanced and sophisticated.
@@
In TokMem, a memory bank stores learned procedures (Figure 1a), where each memory token serves both as an index to a procedure and as a control signal that steers generation, allowing the model to execute targeted tasks with constant-size overhead regardless of the procedure’s complexity. Rather than front-loading procedures as long prompts, TokMem integrates memory tokens directly into the generation process by recalling them on demand at each stage. As illustrated in Figure 1b, TokMem supports multi-step workflows by sequentially recalls procedure tokens such as parsing, searching, and formatting; after the model completes a specific segment of the response, it retrieves the next relevant token to condition the following stage, allowing for modular task composition.
@@
A key advantage of TokMem is that its memory tokens are parameter-isolated from the backbone LLM. Because each procedure is distilled into its own dedicated token, new skills can be added without interfering with or degrading existing ones. This isolation naturally supports continual learning, enabling the model to accumulate procedural skills over time while remaining stable. This mirrors human procedural memory, where skills are gradually acquired through practice and later invoked by contextual cues (Anderson and Lebiere, 1998). Overall, TokMem enables efficient learning and continual expansion of procedural knowledge.
@@
We evaluate TokMem in two complementary settings. First, in atomic recall, we treat each of the 1,000 tasks from Super-Natural Instructions (Wang et al., 2022a) as a distinct procedure. In this setting, TokMem successfully stores and retrieves these procedures efficiently without catastrophic forgetting, outperforming retrieval-based baselines by eliminating the noise and computational overhead of text-based prompts. Second, in compositional recall using function-calling benchmark (Liu et al., 2024b), each tool invocation is treated as an atomic procedure, and answering a query requires chaining multiple such procedures together. Here, TokMem consistently matches or surpasses the performance of parametric fine-tuning while using significantly fewer trainable parameters.
@@
Figure 1: Overview of our TokMem approach. (a) New (colored) memory tokens are interleaved with text sequences and trained via next-token prediction while the LLM backbone remains frozen. (b) At inference time, a query recalls and chains memory tokens (e.g., parse, search, format), enabling multi-step procedural behavior without long prompts.
@@ 2 Method
In this section, we first review the Transformer processing pipeline and the limitations of current textualized prompting (§ 2.1). We then introduce TokMem, detailing how procedures are encoded into a single memory token (§ 2.2), how these tokens are recalled and chained at inference time (§ 2.3), and the renormalization strategy used to stabilize the memory bank (§ 2.4).
@@ 2.1 Textualized Context Engineering
A Transformer (Vaswani et al., 2017) processes a sequence of tokens $(a_{1},\dots,a_{n})\in{\mathbb{N}}^{n}$, where each $a_{i}$ is an integer representing the index of a token (usually a sub-word). The model retrieves the corresponding embedding vector from the embedding layer and produces an input sequence $({\bm{x}}_{1},\dots,{\bm{x}}_{n})^{\top}\in{\mathbb{R}}^{n\times d}$, which is then used to predict the next token in sequence.
Recent advances in prompting can be viewed as textualized context engineering, where the goal is to choose input tokens that steer the model toward desired behavior. For example, chain-of-thought prompting (Wei et al., 2022a) augments the input with intermediate reasoning steps to improve logical inference. Similarly, retrieval-based methods like RAG (Lewis et al., 2020) fetch external documents to ground generation, while recent memory systems such as MemGPT (Packer et al., 2024) retrieve text history into the active context to maintain conversational state. Critically, these approaches all rely on expanding the textual prompt, consuming the limited context window and increasing inference latency due to the quadratic cost of self-attention.
@@ 2.2 TokMem: Procedural Memory as a Token
Our key idea is that frequently reused task procedures can be effectively “compressed” and stored by encoding them into a dedicated memory token, significantly reducing the verbosity of task instructions. Consider $l$ memory tokens. They are added to the vocabulary as special tokens, each represented by an embedding. Thus, they form a memory bank of $l$ special embeddings:
@@
Each ${\bm{m}}_{i}$ is a trainable vector with no direct textual form and represents a unique procedure. For simplicity, we assign each memory token ${\bm{m}}_{i}$ a special token index $a_{m_{i}}\in{\mathbb{N}}$.
To connect these tokens with training, we first describe a single training instance. We define a procedure–response pair, where a procedure is invoked by a memory token ${\bm{m}}_{i}$ and the response is a sequence of textual tokens $(r_{i1},\dots,r_{in})\in{\mathbb{N}}^{n}$ that realizes the procedure in text. For example, the response may specify tool-call arguments or follow a particular output format implied by the procedure.
We formulate training as a supervised task where each procedure–response pair is provided, represented by concatenating a procedural memory token $a_{m_{i}}$ with the embeddings of $(r_{i1},\dots,r_{in})$. Each training instance may contain multiple procedure–response turns, modeling tasks that require multi-step reasoning or composition for a query $q$. Formally, the sequentialized training sequence has the layout
${\bm{a}}=\big(q_{1},\ldots,q_{k},\;\;\underbrace{a_{m_{i}},a_{r_{i1}},a_{r_{i2}},\ldots}_{\text{procedure--response pair}},\;\;\underbrace{a_{m_{j}},a_{r_{j1}},a_{r_{j2}},\ldots}_{\text{procedure--response pair}},\;\;\ldots\big).$ (2)
@@
${\mathcal{L}}({\bm{a}};M)=-\sum_{i>k}\log\Pr(a_{i}\mid{\bm{a}}_{<i};M).$ (3)
The memory embeddings $({\bm{m}}_{1},\dots,{\bm{m}}_{l})$ are shared between the input embedding layer and the LM head. During optimization, these embeddings are trainable, whereas the pre-trained text token embeddings and the backbone LLM remain frozen. Across the training corpus, each memory embedding is exposed to diverse queries and responses, allowing it to learn the underlying procedure representation. We visualize the training process in Figure 1a.
2.3 Inference with Memory Tokens
At inference time, TokMem recalls procedures via memory routing and conditional generation, where routing selects the appropriate memory token for a query. Given $q=(q_{1},\ldots,q_{k})$, the model predicts a distribution over memory tokens from the final hidden state $h_{k}$:
$P(a_{m_{i}}\mid q)\propto\exp\big(\text{logit}(m_{i}\mid h_{k})\big),$ (4)
We select the most probable token $a_{m^{*}}$ and append it to form $q\,;\,a_{m^{*}}]$, after which the model generates the response auto-regressively. For multi-procedure queries, the model can predict additional memory tokens after generating each response segment, enabling token chaining as in Figure 1b. If a query does not match any learned procedure, all memory-token logits may remain low and the model defaults to generating regular text.
In summary, the model decides whether (and how many) memory tokens to recall based on the training layout in (Eq. 2).
@@ 2.4 Stabilizing New Memories
While TokMem supports the modular addition of procedures, integrating new tokens into a frozen backbone presents stability challenges. In continual learning settings, newly optimized memory embeddings are prone to norm inflation (Hou et al., 2019). These new embeddings can dominate the routing logits, suppressing the retrieval of old memories regardless of semantic relevance.
To address this, we introduce renormalization, a lightweight post-update calibration of the memory bank $M\in{\mathbb{R}}^{l\times d}$. Let $A$ and $I$ denote the indices of the new and existing procedural memories, respectively. We estimate the typical scale from the existing set:
$\bar{n}_{I}\;=\;\operatorname{mean}_{j\in I}\bigl\|{\bm{m}}_{j}\bigr\|_{2},$ (5)
and rescale each new embedding as
${\bm{m}}_{i}\;\leftarrow\;{\bm{m}}_{i}\cdot\frac{\bar{n}_{I}}{\|{\bm{m}}_{i}\|_{2}+\varepsilon},\quad i\in A.$ (6)
This operation preserves the directions of newly added embeddings while aligning their magnitudes to the established scale of the memory bank, ensuring smooth integration without overwhelming routing dynamics. The computational overhead is negligible, scaling as $O(|A|d)$.
@@ 3 Experiments
We evaluate TokMem in two complementary scenarios. In atomic memory recall, each task from Super-Natural Instructions (Wang et al., 2022a) is treated as a standalone procedure, where a query maps directly to the desired response. In compositional memory recall, we evaluate multi-procedure tool use on a function-calling dataset (Liu et al., 2024b): each tool invocation is modeled as an atomic procedure, and solving a query requires composing multiple calls. Experiments are conducted on the Qwen (Qwen Team, 2025) and Llama (Llama Team, 2024) families, ranging from 0.5B to 8B.
@@ 3.1 Experimental Setup
Across both settings, we compare TokMem against textualized context engineering, retrieval-augmented memory, and parameter-efficient fine-tuning baselines.
  • Base: In the atomic setting, the model answers queries without demonstrations, providing an empirical lower bound that highlights the need to recall task knowledge.
  • ICL: In the compositional setting, we augment the input with all tool descriptions and prepend two compositional procedure–response demonstrations as a context-engineering baseline.
  • RAG: We use Sentence-BERT (Reimers and Gurevych, 2019) to retrieve relevant demonstrations or tool-use examples and prepend them to the query, following memory-augmented generation (Packer et al., 2024; Chhikara et al., 2025; Xu et al., 2025).
  • Fine-tuning: We fine-tune low-rank adapters (LoRA; Hu et al., 2022), which are inserted into the query and value projections of the transformer, with a parameter count comparable to TokMem. This provides a parametric form of procedural memory.
  • Replay Memory: To mitigate catastrophic forgetting during fine-tuning, we adopt experience replay (Mnih et al., 2015) by maintaining a buffer of previously seen tasks or tools and mixing them into the current training data.
@@
All methods are implemented in HuggingFace Transformers and trained on a single NVIDIA A6000 GPU (48GB) using mixed-precision (bfloat16) training. The backbone models remain frozen: for fine-tuning, only the adapter weights (rank $r=8$) are updated, whereas for TokMem, only the embeddings of newly added procedure IDs are trainable. We expand the tokenizer vocabulary with these procedure IDs and initialize their embeddings by averaging the pretrained embeddings (Hewitt, 2021). For Replay Memory, we mix $20\%$ replayed samples into each batch, using a buffer of $500$ examples refreshed every $10$ tasks in the atomic setting and 1,0001{,}000 examples updated each round in the compositional setting.
We optimize with AdamW using a learning rate of $5\times 10^{-5}$ for fine-tuning and $5\times 10^{-3}$ for TokMem; weight decay is $10^{-2}$ for fine-tuning and zero for TokMem. Training runs for one epoch with batch size $4$ and maximum sequence length $1024$, using teacher forcing and applying the loss only to memory-token and response positions.
Evaluation Metrics.
We evaluate TokMem from two perspectives: (1) memory-token routing accuracy, which measures whether the correct memory tokens are selected, and (2) task performance, which measures task-level generation quality (e.g., ROUGE-L and F1).
Table 1: Atomic recall performance on SNI (ROUGE-L). TokMem consistently outperforms fine-tuning and RAG across models and scales, maintaining strong performance even at 1,000 tasks.
@@
3.2 Atomic Memory Recall
Dataset Details.
We evaluate on Super-Natural Instructions (SNI; Wang et al., 2022a), which provides diverse QA-style natural language tasks. We use ROUGE-L (Lin, 2004) to measure generation quality. Each task is treated as an individual procedure: a query directly invokes the learned procedure to produce the desired response. We sample 1,0001{,}000 English tasks; each task contains $500$ training samples and $50$ test samples. To reflect how memories are typically acquired over time, we introduce tasks sequentially during training rather than all at once, while shuffling examples within each task. We scale the number of tasks from $10$ to 1,0001{,}000 and record checkpoints after training on $\{10,50,200,500,1{,}000\}$ tasks. This resembles incremental domain adaptation (Asghar et al., 2020), where at each checkpoint we evaluate performance across all previously seen tasks. Additional task details are provided in Appendix A.1.
Table 2: Task routing accuracy. TokMem achieves near-perfect routing accuracy at 1,000 tasks, far exceeding the RAG retriever, whose accuracy falls below 80%.